OpenStudy (anonymous):

how can you find mass if you only density given? I know to find mass is density times volume. But im only given density. for volume I have the area already I just need to know how to find velocity when im only given density

OpenStudy (anonymous):

@abb0t help please?

OpenStudy (accessdenied):

Can you post the question where you are having this issue? If you have an area, I suspect you may also have a height or some other length dimension to find volume.

OpenStudy (anonymous):

A 5m diameter spherical balloon contains hydrogen. If the density of the hydrogen is p=.0830 kg/m^3, what is the mass and weight of the hydrogen in the balloon?

OpenStudy (anonymous):

I know how to find weight already I just need help with mass. I know mass is density times volume. I know my density is .0830 I just need help finding volume so I could find mass.

OpenStudy (accessdenied):

They give you a useful piece of information in terms of the shape of the balloon. It is spherical, so its volume would the that of a sphere. Do you remember a formula for that volume, in terms of its radius?

OpenStudy (anonymous):

so the formula that I have is A= 3.14D^2/4 which is my area and I know how to find that too. the formula for V=Av

OpenStudy (anonymous):

All I need help on is finding v which is for velocity.

OpenStudy (accessdenied):

This is the one that I had in mind: \( V = \dfrac{4}{3} \pi r^3 \) Volume in terms of area and velocity doesn't seem right to me. The units would be meters^3 / second, not just volume meters^3.

OpenStudy (anonymous):

It's for an engineering class though.

OpenStudy (accessdenied):

Well, V = A*v would have units [m^2] [m/s] = [m^3/s], which is of volumetric flow rate. That is not simply the volume. So I don't think that formula will be useful here. V = A*h would be a volume, but it applies more for an object with a defined base area A and height h and not a sphere.